Are the following statements true?
 7.5 litres : 15 litres = 5 kg : 10 kg 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

7.5 litres : 15 litres = 5 kg : 10 kg
7.5 / 15 = 1 / 2
5 / 10 = 1 / 2
Hence, True

Thanks